OpenStudy (anonymous):

solve for x: 2/3 (x-2) = 4 - 2/5 - 5/2 2/5 5

hero (hero):

Is it \[\frac{2}{3}(x - 2) = 4\] or \[\frac{2}{3(x - 2)} = 4\]

hero (hero):

It couldn't be the first one because \[\frac{2}{3}(x - 2) = 4\]\[x - 2 = 4\left(\frac{3}{2}\right)\]\[x - 2 = 6\]\[x = 6 + 2\]\[x = 8\] Which isn't any of the answer choices.

hero (hero):

So maybe it's the second one

hero (hero):

\[2 = 4*3(x - 2)\] \[2 = 12(x - 2)\] \[1 = 6(x - 2)\] \[\frac{1}{6} = x - 2\] \[\frac{1}{6} + 2 = x\] \[\frac{1}{6} + \frac{12}{6} = x\] \[\frac{13}{6} = x\] Nope, not that one either. Maybe you have the wrong answer choices listed.

OpenStudy (anonymous):

Ok so im not the only one who got x = 8 lol